OpenStudy (anonymous):

What is the y-intercept of the graph: 3x – 2y = 12 A. (4, 0) B. (4, -6) C. (0, 6) D. (0, -6)

OpenStudy (anonymous):

put x=0 3())-2y=12 -2y=12 y=-6 (0,-6)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw:)